May you please help me with this equation. The pair of values below is from a direct variation. Find the missing number. (2,5) (4,y)
I've taken my first ordered pair (2,5) and set it y=kx which gave me k=5/2.
I'm having trouble figuring the second ordered pair.

Answer by stanbon(75887)

May you please help me with this equation. The pair of values below is from a direct variation. Find the missing number. (2,5) (4,y)
I've taken my first ordered pair (2,5) and set it y=kx which gave me k=5/2.
I'm having trouble figuring the second ordered pair.
-----------
y = kx
5 = k*2
k = 5/2
-----
Equation:
y = (5/2)x
-----
Ans: If x = 4, y = (5/2)4 = 10
-----------------
Cheers,
Stan H.
--------------